OpenStudy (anonymous):

What is the slope and y intercept for 4x+8y>24

OpenStudy (anonymous):

To find y intercept, set x = 0 -4(0)+8y>24 -8y > 24 y > -3

OpenStudy (anonymous):

For slope: Step 1: 4x+8y=-24 Step 2: 8y=-4x-24 Step 3: y=-1/2x-3